write an equation of the line containing the given point and parallel to the given line. Express the answer in y=mx+b form (6,8)x+6y=5. Thanks
Answer by ewatrrr(24785)

 
Hi,
y = mx+b
Line parallel to x + 6y = 5 OR y = (-1/6)x + 5/6 m = (-1/6)
Paralllel lines have identical slopes
New Line:
y = (-1/6)x + b |Using ordered pair (6,8) to solve for b
8 = (-1/6)*6 + b
9 = b
New Line: y = (-1/6)x + 9